A charge of 6.80 pC is spread uniformly throughout the volume of a sphere of radius r = 4.06 cm. What is the magnitude of the electric field at a radial distance of (a)6.17 cm and (b)3.00 cm?
Answer #1

A)

Electric field, E = q/4πε r^2

E = (6.8 x 10^-12) (9 x 10^9)/(0.0617^2) = 16.08 V/m

B)

E = qr/4πε R^3 = (6.8 x 10^-12) (9 x 10^9)x 0.03/0.0406^3

E = 27.43 V/m

Comment in case any doubt please rate my answer....